# Edge-Cloud Routing for Text-to-Image Model with Token-Level Multi-Metric Prediction

### v1

@@ 5 Design of RouteT2I
To achieve the optimization objective in Eq. 2, we propose a text-to-image model routing framework RouteT2I. As shown in Fig. 2, RouteT2I comprises a routing model that predicts the multi-metric generation quality given user prompts and a routing strategy that selects cloud or edge model. In particular, (1) the routing model predicts the quality relationship between images generated on the edge and cloud, namely, Pareto relative superiority. Inspired by the role of prompts in generation processes, where prompt as sequences of tokens interact with images through cross-attention to determine the image content, we treat prompt as a set of tokens in the routing model and select key tokens with influential contextual information for each quality metric. Then, the dominant influence of these tokens is accessed by comparing their positive and negative effects; and (2) the routing strategy aims to maximize quality under cost constraints. The key idea is to use Pareto relative superiority to describe the quality disparity of images generated by the cloud and the edge models. Only prompts that show a notable quality gap when generated in the cloud are routed there, while the rest remain on the cost-effective edge.
Figure 2: Overview of RouteT2I. RouteT2I assembles a pair of off-the-shelf edge and cloud text-to-image models and evaluates them using multi-metric quality across diverse prompts. RouteT2I utilizes Pareto relative superiority between qualities as supervision to train the routing model. Then, to balance quality and cost, the routing strategy determines the most suitable model for each user prompt, choosing between the edge or cloud model.
@@ 5.1 Routing Model Architecture Design with Dual-Gate Token Selection MoE
Our model treats the user prompt as a sequence of tokens, whose influences on the image quality differ due to various weights in cross-attention during generation. To focus on key tokens for each metric, we design a token selection gate, where experts align with quality metrics and actively choose the most relevant tokens to simulate metric attention to different image attributes. Only tokens that significantly impact the quality metrics corresponding to the experts are selected, while those with minimal influences are disregarded, reducing potential interference.
@@ 5.1.2 Dual-Gate MoE
A pair of a positive gate $G^{+}$ and a negative gate $G^{-}$ are introduced for different experts to select tokens that have positive and negative impacts on their corresponding quality metrics, while capturing the weights of their influences. The token selection results are denoted as
@@
To eliminate the redundancy of having separate experts for evaluating both positive and negative aspects of the same quality metric, the structure of the expert is revised to allow an expert to assess both aspects simultaneously. This revision acknowledges that both positive and negative evaluations are inherently part of the same quality metric. Specifically, the original linear layer $W_{i}\in\mathbbm{R}^{d\times h}$ of expert on metric $i$ is split into two projection matrices $P_{i}^{+},P_{i}^{-}\in\mathbbm{R}^{d\times l}$ and one score matrix $S_{i}\in\mathbbm{R}^{l\times h}$. Two projection matrices are specific to positive and negative aspects, projecting tokens into the corresponding representations in an $l$-dimensional low-dimensional space. Since $l\ll h,d$, the parameter size and the computation cost are reduced from $O(hd)$ to $O(l(h+d))$, allowing an expert to efficiently focus on both positive and negative aspects of a metric with a low cost. The shared score matrix then evaluates the impact of these representations on metric $i$. Thus, the impact of $t$-th token $T[t]\in\mathbbm{R}^{d}$ on quality metric $i$ for positive or negative aspect $o$ is described as
@@ 5.2 Routing Strategy Design
Given the Pareto relative superiority predicted by routing model, we propose a routing strategy to effectively route user prompts to suitable models to achieve the trade-off between quality and costs. With a preset routing rate, we can determine the proportion of user prompts sent to the cloud, focusing on those where the cloud model significantly outperforms the edge model in quality. Since Pareto relative superiority represents this quality gap, we can efficiently filter prompts by setting a threshold $\alpha$ on the Pareto relative superiority. In particular, the prompts with a Pareto relative superiority above the threshold are better handled by the edge model for cost efficiency, while those below the threshold achieve superior quality when processed in the cloud. Thus, the optimization objective in Eq. 2 can be expressed as:
@@ 6.1 Experimental Setup
Router Multi-Dimensional Metric $\Delta P(\%)$
@@
Edge Model 0.6251 0.6685 0.6076 0.6537 0.5949 0.5575 0.4680 0.5088 0.4860 0.4690 0.00
Cloud Model 0.6337 0.6847 0.6346 0.6703 0.5930 0.5868 0.5134 0.5199 0.5345 0.4972 80.00
[…]
@@
Table 2: The multi-dimensional quality of images generated by the routed edge and cloud text-to-image model with RouteT2I at the routing rate 50%. The higher the metrics, the better.
Dataset. We take COCO2014 [15], a comprehensive resource for object detection, segmentation, and captioning tasks. We select captions from this dataset to serve as the user prompts $\mathcal{X}$. Given these prompts, we generate images $\mathcal{I}_{e},\mathcal{I}_{c}$ using different open-source text-to-image models.
Models. We take stable diffusion models with varying performances and sizes for cloud and edge usage, include Stable Diffusion 1.5 (SD1.5) [25], Stable Diffusion 2.1 (SD2.1) [25], Stable Diffusion XL (SDXL) [21], Stable Diffusion XL-Refiner (XL-Refiner) [21], and Stable Diffusion 3 (SD3) [8]. We adopt the default hyperparameters in their official documentation for image generation.
Baselines and Settings. We first introduce random routing as a baseline, where prompts are randomly assigned to candidate models, given a specified routing rate to the cloud. Additionally, although there does not exist any previous work on text-to-image generation routing, for a comprehensive comparison, we also reproduce several representative routing methods for LLM by adapting them to our scenarios, including RouteLLM [20] with BERT classifier or matrix factorization, Hybrid LLM [7], and ZOOTER [18]. Unless specified in their paper, the routing models are based on Transformers and take the same hyperparameter settings as in our method. All the routing models are implemented in PyTorch 2.3.1 and trained using Adam optimizer with a learning rate of $52e-$ 5, a weight decay of 00, a batch size of $16$, on a NVIDIA 4090D for about 10 epochs.
@@
For image generation quality, we first fix a routing rate $p$ and introduce the wining rate of the selected models over the cloud model after routing:
$w(\mathcal{I}_{r,p})$ =$=\mathbbm{P}\{\mathcal{Q}(I_{r})\geq\mathcal{Q}(I_{c})\ |\ I_{r},I_{c}\in\mathcal{I}_{r,p},\mathcal{I}_{c}\}$ (16)
=$=\mathbbm{P}\{PRS(I_{r},I_{c})\geq\frac{1}{2}\ |\ I_{r},I_{c}\in\mathcal{I}_{r,p},\mathcal{I}_{c}\},$ (17)
where $\mathcal{I}_{r,p}$ denotes the generated images after routing with the routing rate $p$. To give a clear comparison of the improvement, we introduce the winning rate improvement ratio, which measures the wining rate improvement compared to the random baseline over the improvement achieved by the oracle of the optimal route under ideal conditions:
$\Delta w(p)=\frac{w(\mathcal{I}_{r,p})-w(\mathcal{I}_{b,p})}{w(\mathcal{I}_{o,p})-w(\mathcal{I}_{b,p})},$ (18)
where $\mathcal{I}_{b,p}$ and $\mathcal{I}_{o,p}$ denote the generated images using the random baseline and oracle, respectively. To further quantify the quality improvement, we introduce relative performance improvement, which measures the improvement of the selected model with routing design over the edge model in relation to the improvement achieved by the cloud model:
@@
For cost-quality balance, we introduce the cost saving ratio to measure the reduction of the routing rate to the cloud model, compared to the random baseline at a given relative performance improvement $\Delta P$:
@@ 6.2 Main Results
We present the routing performance of our RouteT2I and baselines using SD3 as the cloud model and SD2.1 as the edge model.
Router Routing Rate ($p$)
40% 50% 60% 70% 80%
@@
Table 3: Winning rate improvement ratios $\Delta w$ (%) at different routing rates.
Wining Rates. Tab. 3 shows the wining rate improvement ratios at various routing rates, and our RouteT2I demonstrates a significant improvement across all routing rates. Specifically, at a 40% routing rate, our improvement reaches 30.60%, exceeding the baselines by at least 3.83% and demonstrating strong consistency with the oracle.
Multi-Metric Quality. Tab. 2 presents multiple image quality metrics at a routing rate of 50%. We can observe that RouteT2I outperforms all the baselines on 6 out of 10 metrics, because our design with a multi-metric optimization objective incorporates each metric simultaneously. The overall relative performance improvement of RouteT2I is highest, improving at least 6% than existing baselines.
Router Relative Performance Improvement ($\Delta P$)
40% 50% 60% 70% 80%
@@
Table 4: Cost saving ratio $\gamma$ (%) compared to random routing at different relative performance improvements.
Cost Efficiency. Tab. 4 presents the cost saving ratios at given relative performance improvement targets. We can see that, at a 40% relative performance improvement, RouteT2I reduces the number of cloud serving calls by up to 71.81% compared to random routing, and reduces at least 5.80% compared to other baselines. In other cases, RouteT2I still reduces over 50% compared to random routing, demonstrating high cost efficiency.
(a) Wining rate improvement ratio.
(b) Relative performance improvement.
Figure 4: Visualization of wining rate improvement ratio $\Delta w$ and relative performance improvement $\Delta P$ when varying routing rates $p$ to the cloud.
Visualization. In Fig. 4, we visualize the wining rate improvement ratio and relative performance improvement. The results indicate that when the routing rate is below approximately 70%, RouteT2I significantly improves quality by effectively assigning the appropriate model for each prompt. However, this advantage diminishes at higher routing rates, because most prompts are processed by the cloud model, limiting the functionality of routing. When the routing rate exceeds 80%, the increase in Fig. 4(b) suggests that RouteT2I approaches the performance of the oracle, confirming this observation.
6.3 Routing Performance on Different T2I Models
Figure 5: Winning rate $w$ between commonly used open-source text-to-image models
Cloud Model Edge Model Routing Rate ($p$)
40% 50% 60% 70% 80%
SD3 SD2.1 30.36 25.81 20.32 18.02 21.94
SD3 XL-Refiner 26.40 21.70 18.83 16.26 21.40
SD3 SDXL 27.10 23.35 20.50 16.24 19.75
SD3 SD1.5 25.16 20.28 17.64 14.13 17.94
SD2.1 XL-Refiner 21.83 18.68 20.99 24.78 28.56
SD2.1 SDXL 19.85 16.61 17.04 19.81 25.54
SD2.1 SD1.5 11.62 9.88 9.75 11.72 13.69
XL-Refiner SDXL 10.88 8.98 9.07 9.82 10.51
XL-Refiner SD1.5 18.52 15.74 18.43 21.30 26.54
SDXL SD1.5 16.52 15.88 19.57 23.98 30.49
Table 5: Winning rate improvement ratio $\Delta w$ (%) of our RouteT2I with different edge and cloud text-to-image models.
To verify the generality of our routing design across different T2I model combinations, Tab. 5 shows the performance of RouteT2I when deploying various T2I models on the cloud and the edge. The quality rankings among these models are depicted in Fig. 5. The results demonstrate that RouteT2I achieves significant improvements, particularly when there is a pronounced quality gap between models, such as with SD3 as the cloud model and others as the edge model. In these cases, RouteT2I consistently achieves over 25% of the oracle’s improvement compared to a random baseline at a 40% routing rate. Notably, even with closely related models like XL-Refiner and SDXL, where XL-Refiner is essentially an SDXL model enhanced with an added refinement stage, RouteT2I still reaches about 10% of the oracle’s improvement. These results highlight RouteT2I’s capability to discern subtle differences in ability between cloud and edge models.
@@ 6.4 Ablation Experiment
Table 6: Ablation experiment of RouteT2I, showing winning rate improvement ratio $\Delta w$ (%) on given routing rates.
Tab. 6 presents the results of ablation study on the multi-metric quality optimation objective, the token selection gate, and the dual-gate MoE in RouteT2I. By utilizing multi-dimensional quality metrics and multiple classification heads, RouteT2I can robustly and comprehensively predict image quality. Omitting them leads to a significant performance drop of approximately 2%. Token selection gates have a significant impact at high routing rates, where quality distribution tends to be sparse. By concentrating on key tokens, RouteT2I can effectively pinpoint essential factors, minimizing interference from unrelated tokens. Meanwhile, the dual-gate MoE performs significantly at moderate routing rates, where tokens with critical quality show no significant difference when generated in the cloud or on the edge. The dual-gate can distinguish subtle variations through the contrast between gates.

### v2

@@ 5 Design of RouteT2I
Figure 4: Overview of RouteT2I. RouteT2I assembles a pair of off-the-shelf edge and cloud text-to-image models and evaluates them using multi-metric image quality. RouteT2I utilizes Pareto relative superiority between qualities as supervision to train the routing model. Then, to balance quality and cost, the routing strategy determines the most suitable model for each user prompt, choosing between the edge or cloud model.
@@
To achieve the optimization objective in Eq. 2, we propose a text-to-image model routing framework RouteT2I. As shown in Fig. 4, RouteT2I comprises a routing model that predicts the multi-metric generation quality given user prompts and a routing strategy that selects cloud or edge model. In particular, (1) the routing model predicts the quality relationship between images generated on the edge and cloud, namely, Pareto relative superiority. Inspired by the role of prompts in generation, where prompt as sequences of tokens interact with images through cross-attention to determine the image content, we treat prompt as a set of tokens in the routing model and select key tokens with influential contextual information for each quality metric. Then, the dominant influence of these tokens is accessed by comparing their positive and negative effects; and (2) the routing strategy aims to maximize quality under cost constraints. The key idea is to use Pareto relative superiority to describe the quality disparity of images generated by the cloud and the edge models. Only prompts that show a notable quality gap when generated in the cloud are routed there, while the rest remain on the cost-effective edge.
@@ 5.1 Routing Model Architecture Design with Dual-Gate Token Selection MoE
Our model treats the user prompt as a sequence of tokens, whose influences on the image quality differ due to their weights in cross-attention during generation. To focus on key tokens for each metric, we design a token selection gate, where experts align with quality metrics and actively choose the most relevant tokens to simulate metric attention to different image attributes. Only tokens that significantly impact the quality metrics corresponding to the experts are selected, while those with minimal influences are disregarded, reducing potential interference.
@@ 5.1.2 Dual-Gate MoE
Both a positive gate $G^{+}$ and a negative gate $G^{-}$ are introduced for different experts to select tokens that have positive and negative impacts on their corresponding quality metrics, while capturing the weights of their influences. The token selection results are denoted as
@@
To eliminate the redundancy of having separate experts for evaluating both positive and negative aspects of the same quality metric, the structure of experts is revised to allow an expert to assess both aspects simultaneously. This revision acknowledges that both positive and negative evaluations are inherently part of the same quality metric. Specifically, the original linear layer $W_{i}\in\mathbbm{R}^{d\times h}$ of expert on metric $i$ is split into two projection matrices $P_{i}^{+},P_{i}^{-}\in\mathbbm{R}^{d\times l}$ and one score matrix $S_{i}\in\mathbbm{R}^{l\times h}$. Two projection matrices are specific to positive and negative aspects, projecting tokens into the corresponding representations in an $l$-dimensional low-dimensional space. Since $l\ll h,d$, the parameter size and the computation cost are reduced from $O(hd)$ to $O(l(h+d))$, allowing an expert to efficiently focus on both positive and negative aspects of a metric with a low cost. The shared score matrix then evaluates the impact of these representations on metric $i$. Thus, the impact of $t$-th token $T[t]\in\mathbbm{R}^{d}$ on quality metric $i$ for positive or negative aspect $o$ is described as
@@ 5.2 Routing Strategy Design
Given the Pareto relative superiority predicted by routing model, we propose a routing strategy to effectively route user prompts to suitable models to achieve the trade-off between quality and costs. With a preset routing rate, we can determine the proportion of user prompts sent to the cloud, focusing on those where the cloud model significantly outperforms the edge model in quality. Since Pareto relative superiority represents this quality gap, we can efficiently filter prompts by setting a threshold $\alpha$ on it. In particular, the prompts with a Pareto relative superiority above the threshold are better handled by the edge model for cost efficiency, while those below the threshold achieve superior quality when processed in the cloud. Thus, the optimization objective in Eq. 2 can be expressed as:
@@ 6.1 Experimental Setup
Router Image Quality Metrics $\Delta P(\%)$
@@
Edge Model 0.6251 0.6685 0.6076 0.6537 0.5949 0.5575 0.4680 0.5088 0.4860 0.4690 -
Cloud Model 0.6337 0.6847 0.6346 0.6703 0.5930 0.5868 0.5134 0.5199 0.5345 0.4972 -
Random 0.6294 0.6766 0.6211 0.6620 0.5939 0.5721 0.4907 0.5144 0.5102 0.4831 40.00
@@
Table 2: The multi-dimensional quality of images generated by text-to-image models, with the router selecting between edge and cloud models for each prompt at a 50% routing rate. The higher the metrics, the better.
Dataset. We take COCO2014 [19], a comprehensive dataset for object detection, segmentation, and captioning tasks. We select its captions as user prompts $\mathcal{X}$. Given these prompts, we generate images $\mathcal{I}_{e},\mathcal{I}_{c}$ using different open-source text-to-image models. Similarly, we use the LAION dataset [31], with results shown in Appendix Sec. F.
Models. We take T2I models with varying performances and sizes for cloud and edge usage, include Stable Diffusion 1.5 (SD1.5) [28], Stable Diffusion 2.1 (SD2.1) [28], Stable Diffusion XL (SDXL) [24], Stable Diffusion XL-Refiner (XL-Refiner) [24], Stable Diffusion 3 (SD3) [9], PixArt-$\alpha$ [5], and Infinity [10]. These models use various text encoders, such as CLIP [25], OpenCLIP [7], and T5 [26], and are based on architectures like diffusion, diffusion Transformer and autoregressive. Details are in Appendix Sec. A.
Baselines and Settings. We first introduce random routing as a baseline, where prompts are randomly assigned to candidate models, given a specified routing rate to the cloud. Additionally, although there does not exist any previous work on text-to-image generation routing, for a comprehensive comparison, we also reproduce several representative routing methods for LLM by adapting them to our scenarios, including RouteLLM [23] with BERT classifier or matrix factorization, Hybrid LLM [8], and ZOOTER [21]. Specifically, we keep their text input processing and quality threshold design, but replace their optimization goal with the average image quality across metrics as they only handle single-objective optimization.
Unless specified in their paper, the routing models are based on Transformers and take the same hyperparameters as in our method. All routers are implemented in PyTorch 2.3.1 and trained using Adam optimizer with a learning rate of 2e-5, a weight decay of 0, a batch size of $16$, on a NVIDIA 4090D for about 10 epochs.
@@
For image generation quality, the absolute performance of routing is constrained by the gap between edge and cloud T2I models. Thus, we introduce a set of relative performance metrics for evaluation. We first fix a routing rate $p$ and introduce the win rate of the selected models over the cloud model after routing:
$w(\mathcal{I}_{r,p})$ $=\mathbbm{P}\{\mathcal{Q}(I_{r})\geq\mathcal{Q}(I_{c})\ |\ I_{r},I_{c}\in\mathcal{I}_{r,p},\mathcal{I}_{c}\},$ (16)
where $\mathcal{I}_{r,p}$ denotes the generated images after routing at a routing rate $p$. To intuitively show the improvement in win rate, we introduce the normalized win rate improvement, which is the increase in win rate achieved by a routing method compared to the baseline, divided by the increase achieved by the optimal oracle routing. A higher value indicates that the routing method is closer to the oracle routing:
$\Delta\overline{w}(p)=\frac{w(\mathcal{I}_{r,p})-w(\mathcal{I}_{b,p})}{w(\mathcal{I}_{o,p})-w(\mathcal{I}_{b,p})},$ (17)
where $\mathcal{I}_{b,p}$ and $\mathcal{I}_{o,p}$ denote the generated images using the random baseline and oracle, respectively. To quantify the improvement of image quality metrics, we introduce relative performance improvement, which measures the proportion of the performance gain achieved through routing compared to the edge model, relative to the total possible improvement brought by fully using the cloud model:
@@
For cost-quality balance, we introduce the cost saving to measure the reduction of the routing rate to the cloud model, compared to the random baseline at a given relative performance improvement $\Delta P$:
@@ 6.2 Main Results
We present the routing performance of our RouteT2I and baselines using SD3 as the cloud model and SD2.1 as the edge model. Human evaluation, cost analysis, visual results, and more T2I pair results are provided in Appendix.
Router pp 40% 50% 60% 70% 80%
@@
Table 3: Normalized win rate improvements $\Delta\overline{w}$ (%) at different routing rates, i.e. the ratio of improvement to oracle improvement.
Win Rates. Tab. 3 shows our RouteT2I demonstrates a significant improvements on normalized win rate improvements across all routing rates. Specifically, at a 40% routing rate, our improvement reaches 30.60%, exceeding the baselines by at least 3.83% and demonstrating strong consistency with the oracle.
Image Quality. Tab. 2 shows RouteT2I outperforms all baselines in 6 of 10 metrics. Our multi-metric optimization balances all metrics simultaneously, while others focus on just a few, often at the cost of others. We also achieve the highest relative performance improvement, reaching 83.97% of cloud model’s gain over edge model, surpassing others by at least 6%. Experiments in Appendix Sec. H show our method may lead to performance gains beyond those of the cloud model itself and is effective for T2I models with various text encoders and architectures.
Router $\Delta P$ 40% 50% 60% 70% 80%
@@
Table 4: Cost saving $\gamma$ (%) under a given relative performance improvements ($\Delta P$) target. $\gamma$ is the ratio of savings to the routing rate required by the random baseline.
Cost Efficiency. Tab. 4 presents the cost savings at given relative performance improvement targets. At a 40% relative performance improvement, RouteT2I reduce cloud service needs by up to 71.81% of the calls required by random routing and by at least 5.80% compared to other baselines. In other cases, RouteT2I still reduces over 50% costs, demonstrating high cost efficiency.
Figure 6: Visualization of (a) Normalized win rate improvement $\Delta\overline{w}$ and (b) relative performance improvement $\Delta P$ at varying routing rates $p$ to the cloud.
Visualization. As shown in Fig. 6, RouteT2I performs well across routing rates, demonstrating its ability to predict quality gaps between T2I models and route effectively.
Visual examples are shown in Appendix Sec. G.
6.3 Routing Performance on More T2I Model Pairs
Cloud-Edge pp 40% 50% 60% 70% 80%
SD3 XL-Refiner 27.42 22.39 18.05 16.85 20.89
SD3 SDXL 29.78 26.81 22.35 16.98 21.30
SD3 SD1.5 26.24 20.88 16.19 14.04 19.22
SD2.1 XL-Refiner 21.62 18.31 20.46 25.30 28.66
SD2.1 SDXL 19.46 16.87 17.27 20.42 23.69
SD2.1 SD1.5 10.77 9.28 10.12 12.22 15.76
XL-Refiner SDXL 13.57 10.15 9.74 11.81 12.59
XL-Refiner SD1.5 19.65 16.98 18.87 22.79 25.57
SDXL SD1.5 18.38 17.01 21.57 24.75 32.49
SD3 PixArt-$\alpha$ 32.26 27.59 24.74 30.73 38.97
PixArt-$\alpha$ SD2.1 25.82 21.26 22.55 28.19 33.76
PixArt-$\alpha$ SDXL 16.71 13.57 12.37 16.35 22.14
PixArt-$\alpha$ XL-Refiner 22.51 19.14 18.18 23.72 28.44
PixArt-$\alpha$ SD1.5 22.38 20.52 19.93 25.40 32.43
SD3 Infinity 22.19 19.97 17.77 19.81 25.39
SD2.1 Infinity 21.98 19.86 22.60 28.45 36.18
PixArt-$\alpha$ Infinity 27.69 23.48 24.12 30.89 38.91
Table 5: Normalized win rate improvement $\Delta\overline{w}$ (%) over oracle routing of our RouteT2I with different cloud-edge model pairs.
To test our method across T2I model pairs, Tab. 5 shows RouteT2I’s performance with various cloud and edge models. Results reveal significant improvements, especially for pairs with large quality gaps, such as SD3 on cloud and others on edge, where RouteT2I achieves over 25% of the oracle routing’s improvement at a 40% routing rate. Even with similar model pairs like XL-Refiner and SDXL, it reaches 10% of the oracle’s improvement, demonstrating its ability to discern subtle differences between T2I models.
@@ 6.4 Ablation Experiment
Table 6: Ablation experiment of RouteT2I, showing normalized win rate improvement $\Delta\overline{w}$ (%) on given routing rates.
Tab. 6 presents ablation results for RouteT2I’s multi-metric quality optimization, token selection gate, and dual-gate MoE. The multi-metric quality optimization allows routers to comprehensively assess the image quality gap, significantly enhancing overall performance. Removing this results in a performance decline across all routing rates, particularly at 40% routing rate, where the drop is approximately 3.23%. The token selection gate prioritizes critical tokens, ensuring routing focuses on key quality factors while reducing interference from irrelevant elements. Its removal notably impacts high routing rates, causing a performance decrease of around 2% at 80% routing rate. The dual-gate MoE effectively distinguishes subtle differences between cloud and edge model outputs, proving most effective at medium routing rates, especially near 50%.